# A fully observer-covariant formulation of the fluid dynamics of simple fluids: derivation, simple examples and a generalized Orr-Sommerfeld equation

### v1

@@ Abstract
We present a formalism to describe the motion of a fluid fully which is fully covariant with respect to arbitrary observers. To achieve fully covariance, we write prognostic equations for quantities that belong to the graded exterior algebra of the cotangent bundle of the manifold occupied by the fluid. With the new formalism, we consider problems of stability, and we derive a generalization of the Orr-Sommerfeld equation that describes the evolution of perturbations relative to an arbitrary observer. The latter is applied to cases where the observer is the Lagrangian observer comoving with the background flow.
@@ 1 Introduction
In the fluid frame the equation for the frame circulation becomes
@@
In the Lagrangian literature, $d\lambda^{f}$ is known as the Cauchy integral (Bennett, 2010). For a barotropic fluid, $-dp/\rho=-dg$ for some $0-$form $g$ and so the Cauchy integral is conserved. In this case, Hodge decomposing the frame circulation
@@
The Lagrangian frame is unique among the Leibnizian frames in that $\lambda^{f}$ is not known a priori, but arises from the dynamics itself. Suppose that a Lagrangian frame determines the solution, that is, determines $\lambda^{f}$. The inversion problem consists in determining the coordinate transformation to a frame which need not be Maxwellian, but whose metric structure is known. That is, given the circulation of the Lagrangian frame $\lambda^{f}$, and a set of coordinates $\hat{x}^{j}$ in a frame in terms of which the metric $\hat{g}_{ij}$ is known, we need to determine the coordinate transformations $\hat{x}^{j}=\hat{x}^{j}(t,x^{1},\ldots,x^{n})$.
@@
$\frac{d(dS^{2})}{dt}=\frac{d}{dt}(\hat{g}_{pq}d\hat{x}^{p}\otimes d\hat{x}^{q})=\frac{d}{dt}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{s}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\hat{g}_{pq}\frac{\partial^{2}\hat{x}^{p}}{\partial x^{i}\partial t}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\hat{g}_{pq}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{j}\partial t}\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{s}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\frac{\partial}{\partial x^{i}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{i}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right.+\\ \left.\frac{\partial}{\partial x^{j}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\right)-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{j}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{i}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{s}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{i}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{j}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{i}}\right)+\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-\frac{\partial\hat{x}^{s}}{\partial t}\left(2\hat{g}_{sq}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{n}}\frac{\partial\hat{x}^{n}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{n}}\frac{\partial\hat{x}^{n}}{\partial x^{j}}\frac{\partial\hat{x}^{q}}{\partial x^{i}}-\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\left(\frac{\partial x^{n}}{\partial\hat{x}^{q}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\frac{1}{2}\hat{g}^{sm}\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{p}}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\frac{1}{2}\hat{g}^{sm}\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{p}}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{j}}\frac{\partial\hat{x}^{q}}{\partial x^{i}}-\right.\right.\\ \left.\left.\frac{1}{2}\hat{g}^{sm}\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\left(\frac{\partial x^{n}}{\partial\hat{x}^{q}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\frac{1}{2}\hat{g}^{sm}\left(\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{p}}+\frac{\partial\hat{g}_{sp}}{\partial\hat{x}^{q}}-\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\right)\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\left(\frac{\partial x^{n}}{\partial\hat{x}^{q}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\hat{\Gamma}^{m}_{pq}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\Gamma^{n}_{ij}\right)dx^{i}\otimes dx^{j}=\widetilde{\nabla_{g}\lambda^{f}},$ (23)
@@
$\frac{\partial g}{\partial t}=\widetilde{\nabla_{g}\lambda^{f}},$ (24)
@@
  • For an incompressible flow in its Lagrangian frame the dynamics reduces to a (deceptively) simple form
    $\frac{\partial g}{\partial t}=\widetilde{\nabla_{g}\lambda^{f}},$ (27a)
@@
For this, we linearize (10) assuming that the frame circulation and the geometry of the Lagrangian background frame are given. The Boussinesq approximation consists in replacing $-dp/\rho-d\mathcal{Z}$ with $-bdz$, where $b=-g(\rho-\rho_{0})/\rho_{0}$ is the buoyancy of the fluid.
@@
Here $\Delta$ is the deRham-Laplace operator accounting for diffusion effects, $\overline{b}$ is the background buoyancy profile, $b$ is the buoyancy perturbation and $\lambda$ is the circulation of the perturbation. Note that the absence of the time derivative of the volume element is due to the fact that, for an incompressible background flow, in Lagrangian coordinates the volume element must be constant in time. Consider now the Hodge decomposition $\lambda=d^{\star}\Psi$, [^3] The closed and harmonic components are assumed zero because we assume the perturbation to be incompressible and the manifold simply connected. where $\Psi$ is a $2$ form. For simplicity, we consider a two-dimensional background flow, and two-dimensional disturbances. Thus, $\Psi=\psi\mathfrak{V}$ for some scalar function $\psi$ and
@@
In this section, we will consider examples of Leibnizian frames (we will use the terms observer and frame interchangeably). The coordinates of an observer within a Galileian frame will be "hatted", while "hat-free" variables will denote the coordinates of an observer in the non-inertial frame.
@@
Relative to the Galileian frame, this Leibnizian frame rotates counter-clockwise (assuming $\omega>0$) around the $\hat{x}^{3}$ axis and measures the axial coordinate relative to a surface, which, in the Galileian frame, may not be stationary. Several types of coordinates can be represented this way, e.g. isopycnic coordinates if the surface is a surface of constant density or the so-called $\sigma-$ coordinates if the surface corresponds to a topographic boundary. To calculate the frame circulation, we begin with the null velocity in the inertial frame, that is, the solution to
@@
where $\zeta_{,j}\equiv\partial\zeta/\partial\hat{x}^{j}$ and $\zeta_{,t}\equiv\partial\zeta/\partial t$. The frame circulation in the Leibnizian frame is given by (we use the fact that the metric in the Galileian frame is diagonal with diagonal $(1,(\hat{x}^{1})^{2},1)$)
@@
while the time derivative of the frame circulation
@@
where $g$ is the gravitational acceleration and $p$ the pressure. Under the same approximations, $\zeta_{,tt}/g\ll 1$, therefore, the time derivative of the frame circulation can be ignored. Finally, as long as $v^{3}\zeta_{,3t}/g\ll 1$, and $\zeta_{,1}\zeta_{,2t}/\omega\simeq\zeta_{,2}\zeta_{,1t}/\omega\ll 1$, the contribution of the heaving isopycnals to the frame vorticity can be neglected. It is only when going beyond the thin layer approximation (e.g., going beyond the hydrostatic approximation, which necessarily implies moving toward suprainertial frequencies) that the heaving isopycnals contribution of the frame vorticity should be included.
@@
(Note that as Galileian components we understand $C=C(x(\hat{x},\hat{y}),y(\hat{x},\hat{y}))$ and similarly for $S$) The frame circulation in the Lagrangian frame can be easily calculated
@@
The frame circulation can expressed as
@@
Stability of Couette flows The stability of Couette flows under infinitesimal perturbations, first considered by (Hopf, 2014), was not definitively settled until much later, when Romanov (1973) proved that the Orr-Sommerfield equation yields stable eigenvalues at all Reynolds numbers (see Drazin and Reid, 2004, and references therein). Here, we show that in the Couette frame, the stability can be proven rather easily, resting fundamentally on the fact that in the Couette frame there is no source of vorticity for the perturbation. For simplicity, we choose a system of units based on shear and viscosity, and we choose the shear $S$ of the Couette flow as a unit of shear and the viscosity $\nu$ of the fluid as unit of viscosity. In these units, lengths have dimension $\sqrt{\nu/S}$. Stability is controlled by specializing the generalized Orr-Sommerfeld equation (Eq. 31) to this particular case. We set $\Omega=S=$const., $\overline{b}=b^{\prime}=0$ and we use the metric derived earlier. we obtain
@@
We consider the evolution of a localized perturbation. We write the potential for the circulation of the perturbation as

### v3

@@ Abstract
We present a formalism to describe the motion of a fluid that is fully covariant with respect to arbitrary observers. To achieve full covariance, we write prognostic equations for quantities that belong to the graded exterior algebra of the cotangent bundle of the manifold occupied by the fluid. In particular, equations are that are fully covariant can be written for a purely Lagrangian observer, for which the fluid velocity (qua section of the tangent bundle) is not a meaningful concept. With the new formalism, we consider problems of stability, and we derive a generalization of the Orr-Sommerfeld equation that describes the evolution of perturbations relative to an arbitrary observer. The latter is applied to cases where the observer is the Lagrangian observer comoving with the background flow.
@@ 1 Introduction
In the fluid frame the equation for the frame lambda becomes
@@
In the Lagrangian literature, $d\lambda^{f}$ is known as the Cauchy integral (Bennett, 2010). For a barotropic fluid, $-dp/\rho=-dg$ for some $0-$form $g$ and so the Cauchy integral is conserved. In this case, Hodge decomposing the frame lambda
@@
The Lagrangian frame is unique among the Leibnizian frames in that $\lambda^{f}$ is not known a priori, but arises from the dynamics itself. Suppose that a Lagrangian frame determines the solution, that is, determines $\lambda^{f}$. The inversion problem consists in determining the coordinate transformation to a frame which need not be Maxwellian, but whose metric structure is known. That is, given the lambda of the Lagrangian frame $\lambda^{f}$, and a set of coordinates $\hat{x}^{j}$ in a frame in terms of which the metric $\hat{g}_{ij}$ is known, we need to determine the coordinate transformations $\hat{x}^{j}=\hat{x}^{j}(t,x^{1},\ldots,x^{n})$.
@@
$\frac{d(dS^{2})}{dt}=\frac{d}{dt}(\hat{g}_{pq}d\hat{x}^{p}\otimes d\hat{x}^{q})=\frac{d}{dt}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{s}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\hat{g}_{pq}\frac{\partial^{2}\hat{x}^{p}}{\partial x^{i}\partial t}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\hat{g}_{pq}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{j}\partial t}\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{s}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\frac{\partial}{\partial x^{i}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{i}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right.+\\ \left.\frac{\partial}{\partial x^{j}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\right)-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{j}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{i}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{s}}{\partial t}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{i}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)-\frac{\partial\hat{x}^{p}}{\partial t}\frac{\partial}{\partial x^{j}}\left(\hat{g}_{pq}\frac{\partial\hat{x}^{q}}{\partial x^{i}}\right)+\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-\frac{\partial\hat{x}^{s}}{\partial t}\left(2\hat{g}_{sq}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{n}}\frac{\partial\hat{x}^{n}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{n}}\frac{\partial\hat{x}^{n}}{\partial x^{j}}\frac{\partial\hat{x}^{q}}{\partial x^{i}}-\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\left(\frac{\partial x^{n}}{\partial\hat{x}^{q}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\frac{1}{2}\hat{g}^{sm}\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{p}}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}+\frac{1}{2}\hat{g}^{sm}\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{p}}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{j}}\frac{\partial\hat{x}^{q}}{\partial x^{i}}-\right.\right.\\ \left.\left.\frac{1}{2}\hat{g}^{sm}\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\left(\frac{\partial x^{n}}{\partial\hat{x}^{q}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\frac{1}{2}\hat{g}^{sm}\left(\frac{\partial\hat{g}_{sq}}{\partial\hat{x}^{p}}+\frac{\partial\hat{g}_{sp}}{\partial\hat{x}^{q}}-\frac{\partial\hat{g}_{pq}}{\partial\hat{x}^{s}}\right)\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\left(\frac{\partial x^{n}}{\partial\hat{x}^{q}}\frac{\partial^{2}\hat{x}^{q}}{\partial x^{i}\partial x^{j}}+\hat{\Gamma}^{m}_{pq}\frac{\partial x^{n}}{\partial\hat{x}^{m}}\frac{\partial\hat{x}^{p}}{\partial x^{i}}\frac{\partial\hat{x}^{q}}{\partial x^{j}}\right)\right)dx^{i}\otimes dx^{j}=\\ \left(\frac{\partial u^{*}_{i}}{\partial x^{j}}+\frac{\partial u^{*}_{j}}{\partial x^{i}}-2u^{*}_{n}\Gamma^{n}_{ij}\right)dx^{i}\otimes dx^{j}=2\widetilde{\nabla_{g}\lambda^{f}},$ (23)
@@
$\frac{\partial g}{\partial t}=2\widetilde{\nabla_{g}\lambda^{f}},$ (24)
@@
  • For an inviscid incompressible flow in its Lagrangian frame the dynamics reduces to a (deceptively) simple form
    $\frac{\partial g}{\partial t}=2\widetilde{\nabla_{g}\lambda^{f}},$ (27a)
@@
For this, we linearize (10) assuming that the frame lambda and the geometry of the Lagrangian background frame are given. The Boussinesq approximation consists in replacing $-dp/\rho-d\mathcal{Z}$ with $-bdz$, where $b=-g(\rho-\rho_{0})/\rho_{0}$ is the buoyancy of the fluid.
@@
Here $\Delta$ is the deRham-Laplace operator accounting for diffusion effects, $\overline{b}$ is the background buoyancy profile, $b$ is the buoyancy perturbation and $\lambda$ is the lambda of the perturbation. Note that the absence of the time derivative of the volume element is due to the fact that, for an incompressible background flow, in Lagrangian coordinates the volume element must be constant in time. Consider now the Hodge decomposition $\lambda=d^{\star}\Psi$, [^3] The closed and harmonic components are assumed zero because we assume the perturbation to be incompressible and the manifold simply connected. where $\Psi$ is a $2$ form. For simplicity, we consider a two-dimensional background flow, and two-dimensional disturbances. Thus, $\Psi=\psi\mathfrak{V}$ for some scalar function $\psi$ and
@@
In this section, we will consider examples of Leibnizian frames (we will use the terms observer and frame interchangeably). The coordinates of an observer within a Galileian frame [^4] This condition can be relaxed to be a Maxwellian frame. will be "hatted", while "hat-free" variables will denote the coordinates of an observer in the non-inertial frame.
@@
Relative to the Galileian frame, this Leibnizian frame rotates counter-clockwise (assuming $\omega>0$) around the $\hat{x}^{3}$ axis and measures the axial coordinate relative to a surface, which, in the Galileian frame, may not be stationary. Several types of coordinates can be represented this way, e.g. isopycnic coordinates if the surface is a surface of constant density or the so-called $\sigma-$ coordinates if the surface corresponds to a topographic boundary. To calculate the frame lambda, we begin with the null velocity in the inertial frame, that is, the solution to
@@
where $\zeta_{,j}\equiv\partial\zeta/\partial\hat{x}^{j}$ and $\zeta_{,t}\equiv\partial\zeta/\partial t$. The frame lambda in the Leibnizian frame is given by (we use the fact that the metric in the Galileian frame is diagonal with diagonal $(1,(\hat{x}^{1})^{2},1)$)
@@
while the time derivative of the frame lambda
@@
where $g$ is the gravitational acceleration and $p$ the pressure. Under the same approximations, $\zeta_{,tt}/g\ll 1$, therefore, the time derivative of the frame lambda can be ignored. Finally, as long as $v^{3}\zeta_{,3t}/g\ll 1$, and $\zeta_{,1}\zeta_{,2t}/\omega\simeq\zeta_{,2}\zeta_{,1t}/\omega\ll 1$, the contribution of the heaving isopycnals to the frame vorticity can be neglected. It is only when going beyond the thin layer approximation (e.g., going beyond the hydrostatic approximation, which necessarily implies moving toward suprainertial frequencies) that the heaving isopycnals contribution of the frame vorticity should be included.
@@
(Note that as Galileian components we understand $C=C(x(\hat{x},\hat{y}),y(\hat{x},\hat{y}))$ and similarly for $S$) The frame lambda in the Lagrangian frame can be easily calculated
@@
The frame lambda can expressed as
@@
Stability of Couette flows The stability of Couette flows under infinitesimal perturbations, first considered by (Hopf, 1914), was not definitively settled until much later, when Romanov (1973) proved that the Orr-Sommerfield equation yields stable eigenvalues at all Reynolds numbers (see Drazin and Reid, 2004, and references therein). Here, we show that in the Couette frame, the stability can be proven rather easily, resting fundamentally on the fact that in the Couette frame there is no source of vorticity for the perturbation. For simplicity, we choose a system of units based on shear and viscosity, and we choose the shear $S$ of the Couette flow as a unit of shear and the viscosity $\nu$ of the fluid as unit of viscosity. In these units, lengths have dimension $\sqrt{\nu/S}$. Stability is controlled by specializing the generalized Orr-Sommerfeld equation (Eq. 31) to this particular case. We set $\Omega=S=$const., $\overline{b}=b^{\prime}=0$ and we use the metric derived earlier. we obtain
@@
We consider the evolution of a localized perturbation. We write the potential for the lambda of the perturbation as